OpenStudy (anonymous):

Anyone know how to go about this problem?! Calculate the 5-unit moving average of the function. Plot the function and its moving average on the same graph, as in Example 4. (You may use graphing technology for the plot, but you should compute the moving average analytically.)

OpenStudy (anonymous):

Although I'm pretty sure I have no idea how to do this problem.. I'm quite confident you didn't give enough information. What was the function?

OpenStudy (anonymous):

whoops yeah sorry the function is f(x) = x^(2/3)

OpenStudy (anonymous):

Ahhh, well... I still have no idea..lol. But at least someone else has a better shot at it now :) Good luck!

OpenStudy (anonymous):

Yeah I hope so :) thanks for the double check!

OpenStudy (anonymous):

Anytime! (:

OpenStudy (baldymcgee6):

similar problem http://nz.answers.yahoo.com/question/index?qid=20080810155317AACeTkQ

OpenStudy (anonymous):

\[\frac{1}{n}\int\limits_{x-n}^{x}f(t)dt\]So,\[\frac{1}{5}\int\limits_{x-5}^{x}t^\frac{2}{3}dt\]and,